OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

x=3

OpenStudy (radar):

From second equation , let x=3+y Now subsstitute for x in first equation 4(3+y) + 2y =12 12 +4y +2y = 12 6y=0 y=0 Now solve for x, x=3+y=3+0=3 x=3, y=0

OpenStudy (radar):

Of course, they didn't ask for value of y. No harm.